OpenStudy (anonymous):

what is 3^4 = 8 in logarithmic form

OpenStudy (phi):

you mean 3^4 = 81 , right ?

OpenStudy (phi):

you could take the log base 3 of both sides \[ \log_3{3^4} = \log_3{81} \] which can be written as \[ 4\log_3{3} = \log_3{81} \] you should know that the log base 3 of 3 is 1 \[ 4\cdot 1 = \log_3{81} \] and \[ \log_3{81} =4\]

OpenStudy (anonymous):

@phi no i attached the pic

OpenStudy (phi):

That (unfortunately) makes no sense 3*3*3*3 does not equal 8 I guess it is a typo. and they mean 81

OpenStudy (anonymous):

so you like e the peppe? you like to sukky sukky fo 5 dolla? oh yah u do boi. u like da p3n1s

OpenStudy (anonymous):

@phi , \[3^4 = 8 \rightarrow FALSE\]

OpenStudy (anonymous):

yes ur correct @phi its not possible :)